The length of the rectangular tennis court at Wimbledon is 2 feet longer than twice the width. If the court's perimeter is 238 feet, what are the court's dimensions?
Answer by jgr45(31)

First, understand what a perimeter is. It is the total distance around the borders of the court. That said, begin by translating to algebraic expressions. This will be a system of two equations:
L = 2W + 2
2L + 2W = 238
First, begin by substituting 2w+2 for L:
2(2w+2) + 2w = 238
4w + 4 + 2w = 238
6w + 4 = 238
6w = 234
w = 39
L = 2(39) + 2
L = 78 + 2
L = 80
So your court dimensions are 80' by 39'.